OpenStudy (anonymous):

my question is: Integral ( sec x + 4 cos x) ^ 2

OpenStudy (zarkon):

multiply it out ... then integrate each term

OpenStudy (anonymous):

Ahaa.. Ok can you solve it, because I didnt able to solve it after multiply it out.

OpenStudy (anonymous):

where are you !!??

OpenStudy (zarkon):

what part are you stuck on?

OpenStudy (anonymous):

after multiply it out.. my brain is closed.

OpenStudy (agreene):

\[\int\limits_{}^{} (\sec(x)+4\cos(x))^2 dx= \int\limits_{}^{}sec^2(x)+8cos^2(x)dx\] \[\int\limits_{}^{} \sec^2(x)dx+8\int\limits_{}^{}\cos^2(x)dx\] \[\tan(x)+4(\sin(x)\cos(x)+x) + C\] unless my brain isnt working right today--which is possible, i'm a bit sick.

OpenStudy (agreene):

which it is, should be tan(x)+8(sinxcosx+x)+C

OpenStudy (anonymous):

ummm.. I will solve it like your solving

OpenStudy (zarkon):

\[\int(\sec(x)+4\cos(x))^2 dx= \int(\sec^2(x)+2\cdot4\cdot\sec(x)\cos(x)+16\cos^2(x))dx\] \[= \int(\sec^2(x)+8+16\cos^2(x))dx=\cdots\]